OpenStudy (anonymous):

5 - (6k + 1) = 2[(5k - 3) - (k - 2)] 5 - 6k - 1 = 10k - 6 - 1k + 2 ? Stuck... Show steps please!

jimthompson5910 (jim_thompson5910):

5 - (6k + 1) = 2[(5k - 3) - (k - 2)] 5 - 6k - 1 = 2(5k - 3) - 2(k - 2) 5 - 6k - 1 = 10k - 6 - 2k + 4 -6k + 4 = 8k - 2 -6k - 8k = -2 - 4 -14k = -6 k = -6/(-14) k = 3/7

jimthompson5910 (jim_thompson5910):

You have the right idea, but you forgot to distribute the 2 across every term on the right side.

OpenStudy (anonymous):

Oh, right! Thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome

razor99 (razor99):

@LynnMichael something is wrong with the guy who gave u the answer

OpenStudy (anonymous):

Ok..

jimthompson5910 (jim_thompson5910):

the answer is just fine, I verified it with multiple sources